OpenStudy (anonymous):

If f(x) = 8 – 3|x + 2|, find f(–6) 20 –16 –4 4

OpenStudy (whpalmer4):

again, just plug in the numbers: \[f(x) = 8 - 3|x+2|\]\[f(-6) = 8-3|-6+2| = 8 - 3|-4| = \] What is \(|-4|\)?

OpenStudy (anonymous):

4?

OpenStudy (whpalmer4):

Right. So replace the \(|-4|\) with \(4\) and complete the arithmetic. What do you get?

OpenStudy (whpalmer4):

\[f(-6) = 8 -3(4) = \]

OpenStudy (anonymous):

16?

OpenStudy (whpalmer4):

\[f(-6) = 8 - 3*4 = \]

OpenStudy (whpalmer4):

You do multiplication and division before addition and subtraction, so that is \[f(-6) = 8 - (3*4) = 8 - 12 = \]

OpenStudy (anonymous):

so -4?

OpenStudy (whpalmer4):

if 8-12 = -4, then that would be the correct answer. Does 8-12 = -4?

OpenStudy (anonymous):

yes :D

OpenStudy (whpalmer4):

Good!